Yes I am a homeschooling mom and I have the answer key to this problem but when my daughter worked it she got a negative. Please explain to me what you do make it positive. Here is the way she worked it. 16-x=2,
-16 -16
--------
-x = -14
The books answer is x = 14 ours is -x = -14

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

.
If you started with:
16-x=2
.
Then subtracting 16 from both sides:
16-x-16=2-16
Simplifying:
-x=-14
Finally, we MULTIPLY both sides by a -1:
(-1) (-x)=(-1)(-14)
x = 14



Answer by Alan3354(69443)

Yes I am a homeschooling mom and I have the answer key to this problem but when my daughter worked it she got a negative. Please explain to me what you do make it positive. Here is the way she worked it. 16-x=2,
-16 -16
--------
-x = -14
The books answer is x = 14 ours is -x = -14
If -x equals -14, then +x equals +14. Multiply both sides by -1.